### v1

# Spin-Orbital Intertwined Topological Superconductivity in a Class of Correlated Noncentrosymmetric Materials

Abstract
In this study, we propose an alternative route to achieving topological superconductivity (TSC). Our approach applies to a new class of correlated noncentrosymmetric materials that host two spin-split Fermi surfaces with identical spin textures due to a spin-orbital intertwined effect. Incorporating multi-orbital repulsive Hubbard interactions, we calculate the superconducting pairings of a minimal two-orbital effective model within a spin-fluctuation-mediated superconductivity framework. We find that, depending on the effective Rashba spin-orbit coupling (RSOC) strength and filling level, the Hubbard interaction can drive the leading pairing symmetry into the $A_{1}(S_{\pm})$, $B_{1}$, $B_{2}$ or $B_{2}(d_{\pm})$ irreducible representations (IRs) of the $C_{4v}$ point group. Notably, the $A_{1}(S_{\pm})$ pairing gives rise to a fully gapped TSC characterized by a $Z_{2}$ invariant, while the $B_{2}(d_{\pm})$ pairing results in a nodal TSC. Our analysis reveals that the fully gapped TSC is predominated by spin-singlet regardless of the presence of the spin-triplet components. This distinguishes our model from noncentrosymmetric materials with conventional Rashba-split band structures, where TSC typically emerges near the van Hove singularity and is primarily driven by $p$-wave or $f$-wave spin-triplet pairing. These features enhances its experimental accessibility, and we discuss potential experimental systems for its realization.
Introduction.— TSCs can host Majorana bound states, which are exotic quasi-particles obeying non-Abelian statistics. This unique feature makes TSCs fundamentally fascinating and potentially valuable for applications in quantum computing[1, 2, 3, 4, 5, 6, 7, 8, 9]. Over the past two decades, many efforts have been devoted to realizing TSCs in various material platforms[10, 11, 12, 13, 14, 15, 16, 17, 18]. The general strategy involves utilizing superconducting proximity effects in either momentum space or real space. The former, also known as connate TSCs, is typically achieved in multiband superconducting systems with topological boundary states, such as Fe(Se,Te), PbTaSe2 etc[13, 14, 15, 16, 17, 18, 19, 20]. The latter approach involves the artificial construction of topological insulator (or semiconductor)-superconductor heterostructures[10, 11, 12]. However, these material platforms pose experimental challenges, requiring high interface quality, a large induced superconducting gap, and precise control of the chemical potential.
Unlike proximity-effect-induced TSCs, intrinsic TSCs exhibit topological properties throughout their entire volume, making them robust against perturbations at boundaries or interfaces[1]. This characteristic provides a significant advantage, positioning intrinsic TSCs as strong candidates for building quantum computing platforms. The pairing symmetry in intrinsic TSCs is typically $p$-wave spin-triplet, which generally arises only in unconventional superconducting systems driven by non-electron-phonon coupling mechanisms, making such materials exceedingly rare in nature. Current candidates include uranium- and cerium-based heavy fermion compounds, such as UTe2, CePt3Si[21, 22, 23, 24, 25]. Consequently, the exploration of new material systems for realizing TSCs remains an important and active area of research.
A comprehensive analysis of material systems exhibiting TSC reveals that their Fermi surfaces typically form robust spin textures, indicating explicit or potential topological characteristics in their electronic structures. Notably, materials such as CePt3Si, Y2C3, A2Cr3As3 (A=K, Rb, Cs)[24, 25, 26, 27, 28, 29], which lack inversion symmetry, display similar features due to RSOC. Comparable spin texture characteristics have also been experimentally observed in the Fermi surfaces of certain copper- and iron-based superconducting systems[30, 31]. These insights suggest that TSCs may emerge in correlated unconventional superconducting systems when RSOC is considered. However, current research on single-orbital Rashba-Hubbard models indicates that the small contribution of the spin-triplet component in these systems is typically insufficient to realize TSCs, except under specific conditions, such as fillings close to the van Hove singularity[32, 33, 34, 35].
To overcome the difficulties in realizing intrinsic TSCs, we propose an alternative approach by introducing a new degree of freedom, such as orbital, layer, or valley. For clarity, we focus on orbital degrees of freedom and consider a two-orbital Rashba-Hubbard model defined on a square lattice. Using the random phase approximation (RPA), we solve the superconducting problem of this model within the framework of spin-fluctuation-mediated superconductivity. We find that as the model parameters vary, the leading superconducting pairing can fall into the $A_{1}(S_{\pm})$, $B_{1}$, $B_{2}$ or $B_{2}(d_{\pm})$ of $C_{4v}$ point group. Interestingly, the $A_{1}(S_{\pm})$ pairing induces a sign-change gap function on the two Fermi surfaces with identical spin textures. This sign-change feature is primarily driven by the spin-singlet-pairing components regardless of the presence of the spin-triplet components. Meanwhile, the $A_{1}(S_{\pm})$ pairing leads to full-gap TSC states characterized by the topological invariant $Z_{2}$. This TSC is extraordinary and different from the ones driven by predominated $p$-wave or $f$-wave spin-triplet pairing. Additionally, the $B_{2}(d_{\pm})$ pairing results in a nodal TSC state with coexistence of zero-energy flat-band and dispersive Majorana states at the system’s boundary.
Figure 1: (a) The two-orbital ($d_{x^{2}-y^{2}},d_{xy}$) unconventional Rashba model defined on a square lattice with $C_{4v}$ symmetry. The relations between the parameters here and those in Eq. 2 are $t_{1/2}=(t_{x^{2}-y^{2}}\pm t_{xy})/2$ and $\lambda_{R}=(\lambda_{R1}+\lambda_{R2})/2$. (b)-(c) The representative band structure and Fermi surfaces for electron doping. The black arrows on the Fermi surfaces denote the effective spin and the color bar labels the renormalized orbital texture $\bra{FS}\tau^{z}\sigma^{0}\ket{FS}_{k}$/$\bra{FS}\tau^{z}\sigma^{0}\ket{FS}_{max}$. The parameters $t_{2}=0$, $\lambda_{R}=0.8$, $\epsilon=1$ and $\mu=-4.5$. (d-e) The calculated bare and RPA longitudinal spin susceptibilities $\chi_{0}^{zz}(\bm{q})$ and $\chi_{RPA}^{zz}(\bm{q})$. The transverse parts are weak[43]. The parameters $U=4.5$,$V=0.94U$ and $J=0.03U$. Other parameters are the same as those in (b) and (c).
Model and method.— The two-orbital unconventional Rashba-Hubbard model is given by
$H=H_{0}+H_{int},$ (1)
with
$H_{0}=\sum_{k}\xi(k)-\lambda_{R}(\tau^{0}+\epsilon\tau^{3})[\bm{\sigma}\times\bm{\nabla_{k}}\gamma(k)]_{z}+\lambda\tau^{2}\sigma^{3}$ (2)
and
$H_{int}$ $=U\sum_{i\alpha}\hat{n}_{i\alpha\uparrow}\hat{n}_{i\alpha\downarrow}+V\sum_{i\alpha<\beta}\sum_{\sigma\sigma^{\prime}}\hat{n}_{i\alpha\sigma}\hat{n}_{i\beta\sigma^{\prime}}$ (3)
$-J\sum_{i,\alpha<\beta}\bm{S}_{i\alpha}\cdot\bm{S}_{i\beta}$
$+J^{\prime}\sum_{i,\alpha<\beta}\sum_{\sigma}c^{\dagger}_{i\alpha\sigma}c^{\dagger}_{i\alpha\bar{\sigma}}c_{i\beta\bar{\sigma}}c_{i\beta\sigma}$
Here, $H_{0}$ can be defined on a lattice with the symmetry of $C_{3v}$, $C_{4v}$ or $C_{6v}$[36, 37, 38]. For clarity and simplicity, we consider square lattice with $C_{4v}$, as shown in Fig. 1 (a). The two orbital are fixed to be $d_{x^{2}-y^{2}}$ and $d_{xy}$. Then, $\xi(k)=2(t_{1}\tau^{0}+t_{2}\tau^{3})\gamma(k)-\mu$ with $\gamma(k)=\cos k_{x}+\cos k_{y}$ and $\mu$ denoting the chemical potential. $\lambda_{R}$ reperents the strength of effective RSOC, and a parameter $\epsilon$ is introduced to measure the anisotropy of two orbitals[39]. $\lambda$ is the strength of on-site SOC. To match the parameter settings in Fig. 1 (a), $t_{1/2}=(t_{x^{2}-y^{2}}\pm t_{xy})/2$ and $\lambda_{R}=(\lambda_{R1}+\lambda_{R2})/2$. In the following, $t_{1}$ is set to $-1$, and other parameters are in unit of $|t_{1}|$. $\lambda$ is set to 3 throughout the calculations. Pauli matrices $(\tau^{0},\bm{\tau})$ and $(\sigma^{0},\bm{\sigma})$ span in orbital and spin spaces, respectively. Then, $H_{0}$ spans in the bais $\Psi_{k}=(c_{x^{2}-y^{2},k,\uparrow},c_{x^{2}-y^{2}k,\downarrow},c_{xy,k,\uparrow},c_{xy,k,\downarrow})^{T}$ with $c_{\alpha k,\sigma}$ being the electron annihilation operator. $U$, $V$, $J$ and $J^{\prime}$ reperents the intra- and inter-orbital on-site repulsive Hubbard interactions, Hund’s coupling, and the pairing hopping term, respectively. $\hat{n}_{i\alpha\sigma}$ and $\bm{S}_{i\alpha}$ denote the spin-$\sigma$ electron density operator and spin operator for electrons in orbital $\alpha$ on $i$ site, respectively. For clarity, we adopt the conventional notations $J=J^{\prime}$ and $V=U-2J$ in the following discussion.
The general bare susceptibility can be expressed as follows,
$\chi^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{q},i\nu_{n})$ (4)
$=-T\sum_{\bm{k},i\omega_{n}}G_{l_{3}\sigma_{3}l_{1}\sigma_{1}}(\bm{k},i\omega_{n})G_{l_{2}\sigma_{2}l_{4}\sigma_{4}}(\bm{k},i\omega_{n}-i\nu_{n}),$
with the Matsubara Green function,
$\hat{G}(\bm{k},i\omega_{n})=[i\omega_{n}-H_{0}(\bm{k})]^{-1}.$ (5)
Here, $l_{1-4}$ are orbital index, and $\sigma_{1-4}$ are spin index. $i\omega_{n}$ and $i\nu_{n}$ fermionic and bosonic Matsubara frequencies, respectively. Using RPA, the dressed susceptibility can be calculated by
$\hat{\chi}^{RPA}(\bm{q},i\nu_{n})=[\hat{1}-\hat{\chi}(\bm{q},i\nu_{n})\hat{W}]^{-1}\bm{\chi}(\bm{q},i\nu_{n}).$ (6)
Here, $\hat{W}$ is the bare interaction matrix with non-zero elements $W^{l_{1}\sigma l_{1}\bar{\sigma}}_{l_{1}\sigma l_{1}\bar{\sigma}}=U$, $W^{l_{1}\sigma l_{1}\sigma}_{l_{1}\bar{\sigma}l_{1}\bar{\sigma}}=-U$, $W^{l_{2}\sigma l_{1}\bar{\sigma}}_{l_{2}\sigma l_{1}\bar{\sigma}}=V$, $W^{l_{1}\sigma l_{1}\sigma}_{l_{2}\bar{\sigma}l_{2}\bar{\sigma}}=-V$, $W^{l_{1}\sigma l_{1}\bar{\sigma}}_{l_{2}\sigma l_{2}\bar{\sigma}}=J$, $W^{l_{1}\sigma l_{2}\sigma}_{l_{1}\bar{\sigma}l_{2}\bar{\sigma}}=-J$, $W^{l_{1}\sigma l_{2}\bar{\sigma}}_{l_{1}\sigma l_{2}\bar{\sigma}}=J^{\prime}$, and $W^{l_{1}\sigma l_{2}\sigma}_{l_{2}\bar{\sigma}l_{1}\bar{\sigma}}=-J^{\prime}$.
According to the theory of fluctuation-mediated superconductivity[40, 41, 42], the effective static superconducting pairing interactions can be calculated by considering bubble and ladder diagrams[43], and are expressed as:
$V^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{k},\bm{k^{\prime}})=W^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}+V_{l}(\bm{k},\bm{k^{\prime}})-V_{b}(\bm{k},\bm{k^{\prime}}),$ (7)
with the ladder term
$V_{l}(\bm{k},\bm{k^{\prime}})=[\hat{W}\hat{\chi}^{RPA}\hat{W}]^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{k}+\bm{k^{\prime}}),$ (8)
and the bubble term
$V_{b}(\bm{k},\bm{k^{\prime}})=[\hat{W}\hat{\chi}^{RPA}\hat{W}]^{l_{1}\sigma_{1}l_{4}\sigma_{4}}_{l_{3}\sigma_{3}l_{2}\sigma_{2}}(\bm{k}-\bm{k^{\prime}}).$ (9)
The possible superconducting pairing can be evaluated through solving the linear Eliashaberg equation as follows,
$\lambda\Delta_{l_{1}\sigma_{1}l_{4}\sigma_{4}}(\bm{k})$ (10)
$=T\sum_{\bm{k^{\prime}}i\omega_{n}}\sum_{l_{2}l_{3}\sigma_{2}\sigma_{3}}V^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{k},\bm{k^{\prime}})F_{l_{3}\sigma_{3}l_{2}\sigma_{2}}(\bm{k^{\prime}},i\omega_{n}),$
with
$F_{l_{3}\sigma_{3}l_{2}\sigma_{2}}(\bm{k^{\prime}},i\omega_{n})$ (11)
$=G_{l_{3}\sigma_{3}l\sigma}(\bm{k^{\prime}},i\omega_{n})\Delta_{l\sigma l^{\prime}\sigma^{\prime}}(\bm{k^{\prime}})G_{l_{2}\sigma_{2}l^{\prime}\sigma^{\prime}}(-\bm{k^{\prime}},-i\omega_{n}).$
Here, $\Delta_{l_{1}\sigma_{1}l_{4}\sigma_{4}}(\bm{k})$ is the superconducting pairing function and the leading superconducting instability can be identified by the largest positive eigen-value of $\lambda$.
Fermionology and spin susceptibility.— Previous studies on the conventional Rashba-Hubbard model have focused on cases where the filling is near the van Hove singularity. This condition induces strong ferromagnetic fluctuations, which, in turn, favor the emergence of leading $p$- or $f$-wave spin-triplet superconducting states[32, 34, 35]. In this work, we relax this restriction and consider more general cases of electron doping beyond the van Hove singularity. Note that the hole doping is the mirror of the electron doping here. A typical band structure and Fermi surfaces under electron doping are shown in the Figs. 1 (b) and (c). Two Fermi surfaces have the same spin textures and apparent orbital textures, as shown in Fig. 1 (c).
The bare and RPA spin susceptibilities corresponding to the Fermi surfaces in Fig. 1 (c) can be calculated with Eqs. 4 and 6. The calculated results indicate that the transverse spin susceptibility $\chi^{+-}(\bm{q})$ is consistently much smaller than the longitudinal spin susceptibility $\chi^{zz}(\bm{q})$[43]. Therefore, we focus on $\chi^{zz}(\bm{q})$ as shown in Figs. 1 (d) and (e), which primarily determines the characteristics of spin fluctuations in the system. In both Figs. 1 (d) and (e), in addition to the four peaks near the $(\pi,\pi)$, eight new peaks emerge, as indicated by the white dashed arrows in these panels. These new peaks correspond to the nesting wave vectors of the two spin-split Fermi surfaces, which are consistent with the black dashed arrows shown in Fig. 1(a). This suggests that these eight new peaks play a crucial role in determining the characteristics of the potential superconducting state, distinguishing it from the single-orbital case. We also calculated the spin susceptibility for other doping cases[43] and found significantly different characteristics compared to the single orbital cases.
Phase diagram and sign-change SC states.— Once superconductivity is induced by antiferromagnetic fluctuations, the pairing symmetry of the superconducting state can be classified according to the IRs of the $C_{4v}$ group. Table 1 lists the possible pairings with their form factors up to the next-nearest neighbor and their corresponding IRs. By solving the linear gap equation in Eq. (10), the leading superconducting pairing can be determined.
Table 1: Possible IRs of superconducting pairings under the constraints of group $C_{4v}$. Here, $g_{k}=\sin k_{x}\sin k_{y}$.
Pairing form $\phi_{g/u}$ / $\boldsymbol{d_{g/u}}$ IRs
$i\phi_{g}\sigma^{2}\tau^{0/3(1)}$ 1, $\cos k_{x}+\cos k_{y}$, $\cos k_{x}\cos k_{y}$ $A_{1}(A_{2})$
$\cos k_{x}-\cos k_{y}$ $B_{1}(B_{2})$
$\sin k_{x}\sin k_{y}$ $B_{2}(B_{1})$
$i\phi_{u}\sigma^{2}\tau^{2}$$\{\sin k_{x},\sin k_{y}\}$ $E$
$i(\bm{d_{u}}\cdot\bm{\sigma})\sigma^{2}\tau^{0/3(1)}$ $(-\sin k_{y},\sin k_{x},0)$ $A_{1}(A_{2})$
$(\sin k_{x},\sin k_{y},0)$ $A_{2}(A_{1})$
$(\sin k_{y},\sin k_{x},0)$ $B_{1}(B_{2})$
$(\sin k_{x},-\sin k_{y},0)$ $B_{2}(B_{1})$
$[(0,0,\sin k_{x}),(0,0,\sin k_{y})]$ $E$
$i(\bm{d_{g}}\cdot\bm{\sigma})\sigma^{2}\tau^{2}$ $(0,0,\cos k_{x}+\cos k_{y})$ $A_{1}$
$(0,0,\cos k_{x}-\cos k_{y})$ $B_{1}$
$(0,0,g_{k})$ $B_{2}$
$\{(\cos k_{x/y},0,0),(0,\cos k_{y/x},0)\}$ $E$
$\{(g_{k},0,0),(0,g_{k},0)\}$
Figure 2: The phase diagrams of superconducting states as functions of some parameters. (a)-(b) $\lambda_{R}$-$\mu$ phase diagram. $U=4.5$ and $J=0.03U$ in (a). $U=5.4$ and $J=0.21U$ in (b). (c)-(d) $U$-$J$ phase diagram. $\lambda_{R}=0.9$ and $\mu=-5.3$ in (c). $\lambda_{R}=0.8$ and $\mu=-4.5$ in (d). In (a)-(d), all other parameters are the same as those in Fig. 1 (b). In panel (a), the positions marked by numbers 1 and 2 correspond to the star-marked positions labeled 1 and 2 in (c) and (d). The notation $S_{\pm}$ and $d_{\pm}$ in $A_{1}$ and $B_{2}$ indicates that the superconducting gap functions on the two spin-split Fermi surfaces have opposite signs along the radial direction.
The phase diagrams of the superconducting state with respect to $\lambda_{R}$-$\mu$ and $U$-$J$ are shown in Figs. 2 and Figs. S4 and S6[43]. Over a wide-range parameter space, we identify three possible leading pairing channels: $A_{1}$, $B_{1}$, and $B_{2}$. Notably, the superconducting states in the $A_{1}$ and $B_{2}$ channels exhibit highly unconventional sign-changing gap structures $A_{1}(S_{\pm})$ and $B_{2}(d_{\pm})$, which indicate that the superconducting gap functions on the two spin-split Fermi surfaces have opposite signs along the radial direction as shown in Fig. 3 (a) and Fig. S9 (a)[43].
Figure 3: (a)-(d) The superconducting gap functions projected onto the corresponding Fermi surfaces. The red and blue colors indicate the opposite signs of superconducting gap functions. (a) and (b) correspond to full and $A_{1,1}=i\sigma^{2}\tau^{0}\sigma^{0}\phi_{1k}$ channels in $A_{1}(S_{\pm})$ IR, respectively. (c) The renormalized gap function $\Delta_{A_{1}}(k_{F},\theta)/max(\Delta_{A_{1}}(k_{F},\theta))$ and $\Delta_{A_{1,1}}(k_{F},\theta)/max(\Delta_{A_{1,1}}(k_{F},\theta))$. (d) The weights of six sub-channels in $A_{1}(S_{\pm})$ IR as function as $t_{2}$. (e)-(f) The fitting gap function $\Delta_{1/2}(k)$ in Eq. 13 along different high-symmetry lines $\Gamma-M$ and $\Gamma-X$. Other parameters in (a)-(d) are the same as those in Fig. 1 (c) and of star labeled 2 in Fig. 2 (a) and (d).
To elucidate the origin of this unconventional sign-change superconducting state, we take the $A_{1}(S_{\pm})$ channel as an example. The gap function of noncentrosymmetric systems can be parameterized as follows[44],
$\hat{\Delta}_{k}=[\hat{\phi}_{k}+\bm{\hat{d}_{k}}\cdot\bm{\sigma}]i\sigma^{2}.$ (12)
In the one-orbital Rashba model, the projection of the gap function onto the two spin-split Fermi surfaces can be expressed in a simplified form, $\Delta_{1/2}(k)=\phi_{k}\pm|\hat{d}_{k}|$ with $\phi_{k}$ as a scalar. Thus, the aforementioned sign-change gap structure only occur when the triplet component is relatively strong and has helical $p$- or $f$-wave spin-triplet pairing, such as when the filling is near the van Hove singularity[32, 34, 35]. Any doping away from the van Hove singularity consistently results in predominant d-wave spin-singlet pairing[33]. However, for the two-orbital model considered here, additional possibilities arise. We project the six sub-channels $i\sigma^{2}$($\tau^{0}\sigma^{0}\phi_{1k}$, $\tau^{3}\sigma^{0}\phi_{2k}$, $\tau^{0}\bm{\sigma}\cdot\bm{d_{3k}}$, $\tau^{3}\bm{\sigma}\cdot\bm{d_{4k}}$, $\tau^{1}\bm{\sigma}\cdot\bm{d_{5k}}$, $\tau^{2}\sigma^{3}d_{6k}$) in the spin-orbital representation, listed in Table 1 onto the two Fermi surfaces. Note that the first two sub-channels belong to spin-singlet while the latter four correspond to spin-triplet. The projection of the first sub-channel $i\sigma^{2}\tau^{0}\sigma^{0}\phi_{1k}$ is shown in Fig. 3 (b), while the others are presented in Fig. S7 [43]. We find that their superconducting gap functions exhibit the same projected distribution on the Fermi surfaces, ensuring consistency with the overall superconducting gap projection, as illustrated in Figs. 3 (a), (b) and (c).
[…]
$\Delta_{1/2}(k)=\sqrt{2(1+\lvert B_{\bm{k}}^{\pm}\lvert^{2})}\phi_{1k},$ (13)
with $B^{\pm}_{\bm{k}}=i(\sqrt{\lambda^{2}+h_{\bm{k}}^{2}\lambda_{R}^{2}}\pm\lambda_{R}h_{\bm{k}})/\lambda$ and $h_{\bm{k}}=\sqrt{\sin^{2}k_{x}+\sin^{2}k_{y}}$. In Fig. 3 (e) and (f), we plot $\Delta_{1/2}(k)$ along the $\Gamma-X$ and $\Gamma-M$ directions, respectively. The sign-change feature of the gap function at the two Fermi momenta is clearly visible. Such sign-change gap structure is consistent with the Fermi surface nesting and $\chi^{zz}(\bm{q})$ in Figs. 1 (c)-(e). Similar analysis can be performed for channel $B_{2}(d_{\pm})$, as shown in Figs. S8-S10[43]. The results indicate that its sign-change feature cannot come from the pure spin-singlet channel and depends on the mixing of spin-singlet and spin-triplet components[43], in contrast to the $A_{1}(S_{\pm})$ pairing.
Topological superconductivity— For noncentrosymmetric systems with time-reversal symmetry, the criteria for identifying TSC require that a pair of spin-split Fermi surfaces enclose an odd number of time-reversal invariant momentum points, and the order parameter must have opposite signs on the two Fermi surfaces[45]. Additionally, the fully gapped superconducting state is preferable. According to this criterion, the Fermi surface in Fig. 1 (c) and the corresponding superconducting gap in Fig. 3 (a) or (b) support TSC.
Figure 4: (a) The Wilson loop calculation for the $A_{1}(S_{\pm})$ superconducting state. (b) The quasi-particle spectrum of $A_{1}(S_{\pm})$ superconducting state. Note that the in-gap edge states are double-degeneracy with opposite chriality. Periodic and open boundary conditions are applied along $x$ and $y$ direction, respectively. In both (a) and (b), the parameters are the same as those in Fig. 3 (a).
For the two-dimensional system with time-reversal symmetry considered here, the Chern number is always zero, and the topological nature of the superconducting state can be characterized by a topological invariant $Z_{2}$. Due to the lack of spatial inversion symmetry, the $Z_{2}$ invariant can be determined by calculating the Wannier centers of the superconducting state’s wave functions[43, 46] and verified through the computation of the edge spectrum[47, 48]. Figs. 4 (a) and (b) show the the Wannier center and edge spectrum, respectively, for the superconducting state with the Fermi surface from Fig. 1 (c) and the gap structure from Fig. 3 (b). It is straightforward to conclude that this superconducting state has a topological invariant $Z_{2}=1$. The superconducting state in $B_{2}(d_{\pm})$ channel has nodes, however, it is also topologically nontrivial[43, 49], and exhibites Majorana flat bands and dispersive Majorana states on edges.
Outlook and conclusion— The TSC state driven by predominant $s$-wave spin-singlet pairing in our study has significant implications. First, TSC states realized through $p$- or $f$-wave spin-triplet pairing are extremely rare in real material systems. In contrast, almost all reported superconductors exhibit $s$-wave spin-singlet pairing, including cuprate, iron-based, and nickelate high-temperature superconductors[50, 51, 52]. Second, the splitting of Fermi surfaces accompanied by spin-texture characteristics due to broken inversion symmetry has been experimentally reported in systems such as cuprate and iron-based superconductors[30, 31]. For instance, spin splitting induced by Rashba SOC could potentially be tuned via electric fields or other external controls. Our proposal offers a viable pathway to realize bulk topological superconductivity in high-temperature superconducting systems without relying on proximity effects. If achieved, this could greatly advance the exploration of topological quantum computing based on topological superconductivity.
In summary, we investigated superconductivity in a new class of correlated non-centrosymmetric systems. Using the theory of antiferromagnetic-fluctuation-mediated superconductivity, we identified four possible superconducting states $A_{1}(S_{\pm})$, $B_{1}$, $B_{2}$ and $B_{2}(d_{\pm})$ under different model parameters. all of which are predominantly driven by spin-singlet pairing with $A_{1}$, $B_{1}$ and $B_{2}$ pairing symmetries. Notably, the superconducting state with $A_{1}(S_{\pm})$ IR is predominantly driven by spin-singlet pairing and belongs to a bulk topological superconducting state characterized by a $Z_{2}=1$ topological invariant. Our proposal is different from all present scenarios based on proximity effects and strong $p$- or $f$-wave spin-triplet pairing. We further highlighted the potential realization of such topological superconducting states in certain high-temperature superconductors, providing a new pathway for exploring next-generation high-temperature topological superconductors.
Acknowledgements.
This work was financially supported by the National Key R&D Program of China (Grants No. 2022YFA1403200, No. 2024YFA1613200 and Grant No. 2023YFA1407300), National Natural Science Foundation of China (Grants No. 92265104, No. 12022413 and No. 12447103), the Basic Research Program of the Chinese Academy of Sciences Based on Major Scientific Infrastructures (Grant No. JZHKYPT-2021-08), the CASHIPS Director’s Fund (Grant No. BJPY2023A09), Anhui Provincial Major S&T Project(s202305a12020005), and the High Magnetic Field Laboratory of Anhui Province under Contract No. AHHM-FX-2020-02.

### v2

@@
 Abstract
-In this study, we propose an alternative route to achieving topological superconductivity (TSC).
-Our approach applies to a new class of correlated noncentrosymmetric materials that host two spin-split Fermi surfaces with identical spin textures due to a spin-orbital intertwined effect.
-Incorporating multi-orbital repulsive Hubbard interactions, we calculate the superconducting pairings of a minimal two-orbital effective model within a spin-fluctuation-mediated superconductivity framework.
-We find that, depending on the effective Rashba spin-orbit coupling (RSOC) strength and filling level, the Hubbard interaction can drive the leading pairing symmetry into the $A_{1}(S_{\pm})$, $B_{1}$, $B_{2}$ or $B_{2}(d_{\pm})$ irreducible representations (IRs) of the $C_{4v}$ point group.
-Notably, the $A_{1}(S_{\pm})$ pairing gives rise to a fully gapped TSC characterized by a $Z_{2}$ invariant, while the $B_{2}(d_{\pm})$ pairing results in a nodal TSC.
-Our analysis reveals that the fully gapped TSC is predominated by spin-singlet regardless of the presence of the spin-triplet components.
-This distinguishes our model from noncentrosymmetric materials with conventional Rashba-split band structures, where TSC typically emerges near the van Hove singularity and is primarily driven by $p$-wave or $f$-wave spin-triplet pairing.
-These features enhances its experimental accessibility, and we discuss potential experimental systems for its realization.
-Introduction.— TSCs can host Majorana bound states, which are exotic quasi-particles obeying non-Abelian statistics.
+In this study, we propose an alternative route to realizing topological superconductivity (TSC).
+Our approach targets a new class of correlated noncentrosymmetric materials that host two spin-split Fermi surfaces due to spin-orbital intertwined effects.
+By investigating the superconducting pairings based on two-orbital Rashba-Hubbard model on a square lattice within a spin-fluctuation-mediated pairing framework, we find that, depending on model parameters, the leading superconducting state belongs to the $A_{1}(S_{\pm})$, $B_{2}$ or $B_{2}(d_{\pm})$ irreducible representations (IRs) of the $C_{4v}$ point group.
+Notably, the $A_{1}(S_{\pm})$ state features a sign-changed gap between the two spin-split Fermi surfaces and exhibits two key characteristics: (i) it is a parity-mixed state, with the ratio of even-parity to odd-parity components tunable via the Rashba spin-orbit coupling (RSOC) strength and onsite orbital-dependent potential; (ii) it is a fully gapped TSC, characterized by a $Z_{2}$ topological invariant.
+When RSOC is strong, the $B_{2}(d_{\pm})$ state can emerge and exhibits both parity-mixed and nodal TSC features.
+Further analysis reveals that the fully gapped TSC can be predominated by spin-singlet, despite of the presence of the spin-triplet components.
+This distinguishes our case from single-orbital noncentrosymmetric materials, where TSC arises from $p$-wave or $f$-wave spin-triplet pairing.
+In those systems, the Fermi level must be tuned near a van Hove singularity to enhance ferromagnetic fluctuations that drive the spin-triplet pairing.
+These distinctions enhance the experimental feasibility of our model.
+I Introduction
+TSCs can host Majorana bound states, which are exotic quasi-particles obeying non-Abelian statistics.
 This unique feature makes TSCs fundamentally fascinating and potentially valuable for applications in quantum computing[1, 2, 3, 4, 5, 6, 7, 8, 9].
@@
 A comprehensive analysis of material systems exhibiting TSC reveals that their Fermi surfaces typically form robust spin textures, indicating explicit or potential topological characteristics in their electronic structures.
-Notably, materials such as CePt3Si, Y2C3, A2Cr3As3 (A=K, Rb, Cs)[24, 25, 26, 27, 28, 29], which lack inversion symmetry, display similar features due to RSOC.
-Comparable spin texture characteristics have also been experimentally observed in the Fermi surfaces of certain copper- and iron-based superconducting systems[30, 31].
+Notably, materials such as CePt3Si, Y2C3, A2Cr3As3 (A=K, Rb, Cs)[24, 25, 26, 27, 28, 29], which lack central inversion symmetry, display similar features due to RSOC.
+Comparable spin texture characteristics have also been experimentally observed in the Fermi surfaces of certain copper- and iron-based superconductors[30, 31].
 These insights suggest that TSCs may emerge in correlated unconventional superconducting systems when RSOC is considered.
 However, current research on single-orbital Rashba-Hubbard models indicates that the small contribution of the spin-triplet component in these systems is typically insufficient to realize TSCs, except under specific conditions, such as fillings close to the van Hove singularity[32, 33, 34, 35].
-To overcome the difficulties in realizing intrinsic TSCs, we propose an alternative approach by introducing a new degree of freedom, such as orbital, layer, or valley.
-For clarity, we focus on orbital degrees of freedom and consider a two-orbital Rashba-Hubbard model defined on a square lattice.
+To address the challenges in realizing intrinsic TSCs, we propose an alternative approach by introducing a new degree of freedom, such as orbital, layer, or valley.
+For clarity, we focus on orbital degrees of freedom and consider a two-orbital Rashba model, in which the two spin-split Fermi surfaces can exhibit either conventional or unconventional spin textures.
+This transition can be tuned by an on-site orbital-dependent potential together with onsite SOC, reflecting the evolution from the single-orbital limit to the two-orbital limit.
 Using the random phase approximation (RPA), we solve the superconducting problem of this model within the framework of spin-fluctuation-mediated superconductivity.
-We find that as the model parameters vary, the leading superconducting pairing can fall into the $A_{1}(S_{\pm})$, $B_{1}$, $B_{2}$ or $B_{2}(d_{\pm})$ of $C_{4v}$ point group.
-Interestingly, the $A_{1}(S_{\pm})$ pairing induces a sign-change gap function on the two Fermi surfaces with identical spin textures.
-This sign-change feature is primarily driven by the spin-singlet-pairing components regardless of the presence of the spin-triplet components.
-Meanwhile, the $A_{1}(S_{\pm})$ pairing leads to full-gap TSC states characterized by the topological invariant $Z_{2}$.
-This TSC is extraordinary and different from the ones driven by predominated $p$-wave or $f$-wave spin-triplet pairing.
-Additionally, the $B_{2}(d_{\pm})$ pairing results in a nodal TSC state with coexistence of zero-energy flat-band and dispersive Majorana states at the system’s boundary.
-Figure 1: (a) The two-orbital ($d_{x^{2}-y^{2}},d_{xy}$) unconventional Rashba model defined on a square lattice with $C_{4v}$ symmetry.
-The relations between the parameters here and those in Eq.
-2 are $t_{1/2}=(t_{x^{2}-y^{2}}\pm t_{xy})/2$ and $\lambda_{R}=(\lambda_{R1}+\lambda_{R2})/2$.
-(b)-(c) The representative band structure and Fermi surfaces for electron doping.
-The black arrows on the Fermi surfaces denote the effective spin and the color bar labels the renormalized orbital texture $\bra{FS}\tau^{z}\sigma^{0}\ket{FS}_{k}$/$\bra{FS}\tau^{z}\sigma^{0}\ket{FS}_{max}$.
-The parameters $t_{2}=0$, $\lambda_{R}=0.8$, $\epsilon=1$ and $\mu=-4.5$.
-(d-e) The calculated bare and RPA longitudinal spin susceptibilities $\chi_{0}^{zz}(\bm{q})$ and $\chi_{RPA}^{zz}(\bm{q})$.
-The transverse parts are weak[43].
-The parameters $U=4.5$,$V=0.94U$ and $J=0.03U$.
-Other parameters are the same as those in (b) and (c).
-Model and method.— The two-orbital unconventional Rashba-Hubbard model is given by
+We find that, depending on the strength of onsite SOC and effective RSOC, onsite orbital-dependent potential and filling level, the Hubbard interaction can drive the leading pairing symmetry into the $A_{1}(S_{\pm})$, $B_{2}$ or $B_{2}(d_{\pm})$ IRs of the $C_{4v}$ point group.
+Notably, the $A_{1}(S_{\pm})$ IR is conformed to host an $S_{\pm}$-wave pairing state, which lead to a fully gapped TSC characterized by a $Z_{2}$ topological invariant.
+Furthermore, the $S_{\pm}$-wave TSC is a parity-mixed state, with the parity-mixed ratio tunable via the strength of onsite SOC and effective RSOC and onsite orbital-dependent potential.
+This tunability offers a promising platform for exploring parity-mixed superconducting states.
+More significantly, the TSC can be predominatntly driven by spin-singlet component, despite of the spin-triplet contribution.
+We also find the $B_{2}(d_{\pm})$ state can emerge under the strong RSOC and exhibits both parity-mixed and nodal TSC features.
+II Model and method
+The two-orbital Rashba-Hubbard model[36, 37, 38, 39] is given by
 $H=H_{0}+H_{int},$ (1)
 with
-$H_{0}=\sum_{k}\xi(k)-\lambda_{R}(\tau^{0}+\epsilon\tau^{3})[\bm{\sigma}\times\bm{\nabla_{k}}\gamma(k)]_{z}+\lambda\tau^{2}\sigma^{3}$ (2)
-and
+$H_{0}=\sum_{k}\xi(k)-\lambda_{R}(\tau^{0}+\epsilon\tau^{3})[\bm{\sigma}\times\bm{\nabla_{k}}\gamma_{1}(k)]_{z}+\lambda_{so}\tau^{2}\sigma^{3},$ (2)
+in which, the SOC terms have the similar forms of those in the Kane-Mele model [40].RSOC is origined from $H_{RSOC}\propto(\nabla V\times p)\cdot\sigma$ under the condition $V(z)\neq V(-z)$ with $V(\pm z)$being the polarization fields along the ploar axis defined in Fig.
+1 (a).
+Here, $H_{0}$ can be defined on a square lattice with $C_{4v}$ symmetry, as shown in Fig.
+1 (a) .
+The two orbitals are fixed to be $d_{x^{2}-y^{2}}$ and $d_{xy}$.
+Then, $\xi(k)=2(t^{nn}_{1}\tau^{0}+t^{nn}_{2}\tau^{3})\gamma_{1}(k)+4(t^{nnn}_{1}\tau^{0}+t^{nnn}_{2}\tau^{3})\gamma_{2}(k)-\epsilon_{0}^{onsite}\tau^{3}-\mu$ with $\gamma_{1}(k)=\cos k_{x}+\cos k_{y}$, $\gamma_{2}(k)=\cos k_{x}\cos k_{y}$, $\epsilon_{0}^{onsite}$ indicating the difference of on-site orbital-dependent potential and $\mu$ denoting the chemical potential.
+$\lambda_{R}$ represents the strength of effective RSOC, and a parameter $\epsilon$ is introduced to measure the RSOC anisotropy of two orbitals[41].
+$\lambda_{so}$ is the strength of on-site SOC.
+To match the parameter settings in Fig.
+1 (a), $t^{nn/nnn}_{1/2}=(t^{nn/nnn}_{x^{2}-y^{2}}\pm t^{nn/nnn}_{xy})/2$ with $t^{nn/nnn}_{x^{2}-y^{2}/xy}$ denoting the nearest (nn) and next-nearest (nnn) neighbor hoppings for two $d_{x^{2}-y^{2}/xy}$ orbitals and $\lambda_{R}=(\lambda_{R1}+\lambda_{R2})/2$.
+In the following, $t^{nn}_{1}$ is set to 1-1, and other parameters are in unit of $|t^{nn}_{1}|$.
+$\lambda_{so}$ is set to $3+\epsilon_{0}^{onsite}$ throughout the calculations.
+Pauli matrices $(\tau^{0},\bm{\tau})$ and $(\sigma^{0},\bm{\sigma})$ span in orbital and spin spaces, respectively.
+Then, $H_{0}$ spans in the basis $\Psi_{k}=(c_{x^{2}-y^{2},k,\uparrow},c_{x^{2}-y^{2}k,\downarrow},c_{xy,k,\uparrow},c_{xy,k,\downarrow})^{T}$ with $c_{\alpha k,\sigma}$ being the electron annihilation operator.
+The Hamiltonian $H_{0}$ in Eq.
+(2) yields the representative band structure, Fermi surfaces, and spin textures shown in Fig.
+1.
+We find that when the parameter $\epsilon_{0}^{onsite}<-0.46$ or $\lambda_{so}<2.54$, the two spin-split Fermi surfaces display a conventional spin texture, indicating that the model reduces to the single-orbital limit.
+In contrast, when $\epsilon_{0}^{onsite}>-0.46$ or $\lambda_{so}>2.54$, the two spin-split Fermi surfaces exhibit an unconventional spin texture, signaling that the model is in the two-orbital limit.
+This property will impact the magnitude of transverse susceptibility and ultimately affect the effective inter-pocket interaction.
+The below discussions will cover both limits.
+Figure 1: (a) The two-orbital ($d_{x^{2}-y^{2}},d_{xy}$) unconventional Rashba model defined on a square lattice with $C_{4v}$ symmetry.
+The polar axis is defined as the axis perpendicular to the square lattice plane, indicating that along the two opposite directions of this axis the system experiences different polarization fields $V(z)\neq V(-z)$.
+This asymmetry can in turn generate Rashba-type SOC.
+(b) The transition from single-orbital limit with conventional spin texture (left part in (b)) to two-orbital limit with unconventional spin texture (right part in (b)) controlled by $\epsilon_{0}^{onsite}$ or $\lambda_{so}$.
+The circles with cyan color denote the Fermi surfaces and the red arrows indicate the spin texture.
+(c) The band structures under different values of parameters $\epsilon_{0}^{onsite}$ or $\lambda_{so}$.
+From left to right, the parameters are set to be $\lambda_{so}=0$, $\lambda_{so}=1.5$ and $\lambda_{so}=3$ respectively.
+$\lambda_{so}=3+\epsilon_{0}^{onsite}$ is also fixed in order to prevent significant changes in the band structure.
+Other parameters are set as follows: $t^{nn}_{2}=-0.1$, $\lambda_{R}=0.5$, $\epsilon=1$, $t^{nnn}_{1}=0.0125$, $t^{nnn}_{2}=0.0375$, the occupied particle number per site is set to be $n=0.5$.
+Here, the coordinates of high-symmetry points are defined as follows: $\Gamma=(0,0)$; $X=(\pm\pi,0)$, $(0,\pm\pi)$; $M=(\pm\pi,\pm\pi)$.
+The two-orbital Hubbard interactions can be expressed as:
 $H_{int}$ $=U\sum_{i\alpha}\hat{n}_{i\alpha\uparrow}\hat{n}_{i\alpha\downarrow}+V\sum_{i\alpha<\beta}\sum_{\sigma\sigma^{\prime}}\hat{n}_{i\alpha\sigma}\hat{n}_{i\beta\sigma^{\prime}}$ (3)
 $-J\sum_{i,\alpha<\beta}\bm{S}_{i\alpha}\cdot\bm{S}_{i\beta}$
-$+J^{\prime}\sum_{i,\alpha<\beta}\sum_{\sigma}c^{\dagger}_{i\alpha\sigma}c^{\dagger}_{i\alpha\bar{\sigma}}c_{i\beta\bar{\sigma}}c_{i\beta\sigma}$
-Here, $H_{0}$ can be defined on a lattice with the symmetry of $C_{3v}$, $C_{4v}$ or $C_{6v}$[36, 37, 38].
-For clarity and simplicity, we consider square lattice with $C_{4v}$, as shown in Fig.
-1 (a).
-The two orbital are fixed to be $d_{x^{2}-y^{2}}$ and $d_{xy}$.
-Then, $\xi(k)=2(t_{1}\tau^{0}+t_{2}\tau^{3})\gamma(k)-\mu$ with $\gamma(k)=\cos k_{x}+\cos k_{y}$ and $\mu$ denoting the chemical potential.
-$\lambda_{R}$ reperents the strength of effective RSOC, and a parameter $\epsilon$ is introduced to measure the anisotropy of two orbitals[39].
-$\lambda$ is the strength of on-site SOC.
-To match the parameter settings in Fig.
-1 (a), $t_{1/2}=(t_{x^{2}-y^{2}}\pm t_{xy})/2$ and $\lambda_{R}=(\lambda_{R1}+\lambda_{R2})/2$.
-In the following, $t_{1}$ is set to $-1$, and other parameters are in unit of $|t_{1}|$.
-$\lambda$ is set to 3 throughout the calculations.
-Pauli matrices $(\tau^{0},\bm{\tau})$ and $(\sigma^{0},\bm{\sigma})$ span in orbital and spin spaces, respectively.
-Then, $H_{0}$ spans in the bais $\Psi_{k}=(c_{x^{2}-y^{2},k,\uparrow},c_{x^{2}-y^{2}k,\downarrow},c_{xy,k,\uparrow},c_{xy,k,\downarrow})^{T}$ with $c_{\alpha k,\sigma}$ being the electron annihilation operator.
-$U$, $V$, $J$ and $J^{\prime}$ reperents the intra- and inter-orbital on-site repulsive Hubbard interactions, Hund’s coupling, and the pairing hopping term, respectively.
+$+J^{\prime}\sum_{i,\alpha<\beta}\sum_{\sigma}c^{\dagger}_{i\alpha\sigma}c^{\dagger}_{i\alpha\bar{\sigma}}c_{i\beta\bar{\sigma}}c_{i\beta\sigma}.$
+Here, $U$, $V$, $J$ and $J^{\prime}$ represent the intra- and inter-orbital on-site repulsive Hubbard interactions, Hund’s coupling, and the pairing hopping term, respectively.
 $\hat{n}_{i\alpha\sigma}$ and $\bm{S}_{i\alpha}$ denote the spin-$\sigma$ electron density operator and spin operator for electrons in orbital $\alpha$ on $i$ site, respectively.
-For clarity, we adopt the conventional notations $J=J^{\prime}$ and $V=U-2J$ in the following discussion.
+For clarity, we adopt the conventional notations $J=J^{\prime}$ and V=U2JV=U-2J in the following discussion.
+[…]
 The general bare susceptibility can be expressed as follows,
@@
 Here, $\hat{W}$ is the bare interaction matrix with non-zero elements $W^{l_{1}\sigma l_{1}\bar{\sigma}}_{l_{1}\sigma l_{1}\bar{\sigma}}=U$, $W^{l_{1}\sigma l_{1}\sigma}_{l_{1}\bar{\sigma}l_{1}\bar{\sigma}}=-U$, $W^{l_{2}\sigma l_{1}\bar{\sigma}}_{l_{2}\sigma l_{1}\bar{\sigma}}=V$, $W^{l_{1}\sigma l_{1}\sigma}_{l_{2}\bar{\sigma}l_{2}\bar{\sigma}}=-V$, $W^{l_{1}\sigma l_{1}\bar{\sigma}}_{l_{2}\sigma l_{2}\bar{\sigma}}=J$, $W^{l_{1}\sigma l_{2}\sigma}_{l_{1}\bar{\sigma}l_{2}\bar{\sigma}}=-J$, $W^{l_{1}\sigma l_{2}\bar{\sigma}}_{l_{1}\sigma l_{2}\bar{\sigma}}=J^{\prime}$, and $W^{l_{1}\sigma l_{2}\sigma}_{l_{2}\bar{\sigma}l_{1}\bar{\sigma}}=-J^{\prime}$.
-According to the theory of fluctuation-mediated superconductivity[40, 41, 42], the effective static superconducting pairing interactions can be calculated by considering bubble and ladder diagrams[43], and are expressed as:
+According to the theory of fluctuation-mediated superconductivity[42, 43, 44], the effective static superconducting pairing interactions can be calculated by considering bubble and ladder diagrams shown in Fig.
+A1, and are expressed as:
 $V^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{k},\bm{k^{\prime}})=W^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}+V_{l}(\bm{k},\bm{k^{\prime}})-V_{b}(\bm{k},\bm{k^{\prime}}),$ (7)
@@
 Here, $\Delta_{l_{1}\sigma_{1}l_{4}\sigma_{4}}(\bm{k})$ is the superconducting pairing function and the leading superconducting instability can be identified by the largest positive eigen-value of $\lambda$.
-Fermionology and spin susceptibility.— Previous studies on the conventional Rashba-Hubbard model have focused on cases where the filling is near the van Hove singularity.
+More details can be found in Appdenix A.
+III Fermionology and spin susceptibility
+Previous studies on the conventional Rashba-Hubbard model have focused on cases where the filling is near the van Hove singularity.
 This condition induces strong ferromagnetic fluctuations, which, in turn, favor the emergence of leading $p$- or $f$-wave spin-triplet superconducting states[32, 34, 35].
 In this work, we relax this restriction and consider more general cases of electron doping beyond the van Hove singularity.
-Note that the hole doping is the mirror of the electron doping here.
 A typical band structure and Fermi surfaces under electron doping are shown in the Figs.
-1 (b) and (c).
-Two Fermi surfaces have the same spin textures and apparent orbital textures, as shown in Fig.
-1 (c).
-The bare and RPA spin susceptibilities corresponding to the Fermi surfaces in Fig.
-1 (c) can be calculated with Eqs.
-4 and 6.
-The calculated results indicate that the transverse spin susceptibility $\chi^{+-}(\bm{q})$ is consistently much smaller than the longitudinal spin susceptibility $\chi^{zz}(\bm{q})$[43].
-Therefore, we focus on $\chi^{zz}(\bm{q})$ as shown in Figs.
-1 (d) and (e), which primarily determines the characteristics of spin fluctuations in the system.
-In both Figs.
-1 (d) and (e), in addition to the four peaks near the $(\pi,\pi)$, eight new peaks emerge, as indicated by the white dashed arrows in these panels.
-These new peaks correspond to the nesting wave vectors of the two spin-split Fermi surfaces, which are consistent with the black dashed arrows shown in Fig.
-1(a).
-This suggests that these eight new peaks play a crucial role in determining the characteristics of the potential superconducting state, distinguishing it from the single-orbital case.
-We also calculated the spin susceptibility for other doping cases[43] and found significantly different characteristics compared to the single orbital cases.
-Phase diagram and sign-change SC states.— Once superconductivity is induced by antiferromagnetic fluctuations, the pairing symmetry of the superconducting state can be classified according to the IRs of the $C_{4v}$ group.
+1 (c) and Figs.
+2(a), (f) respectively.
+To analyze the behavior of the spin susceptibilities, we define the scalar forms of longitudinal and transverse spin susceptibilities $\chi^{RPA}_{zz}(\bm{q})$ and $\chi^{RPA}_{+-}(\bm{q})$ as follows,
+$\chi^{RPA}_{zz}(\bm{q})=\sum_{l_{1},l_{2}}\sum_{\{\sigma\}}\sigma^{z}_{\sigma_{1}\sigma_{2}}\sigma^{z}_{\sigma_{4}\sigma_{3}}\chi^{l_{1}\sigma_{1}l_{1}\sigma_{2}}_{l_{2}\sigma_{3}l_{2}\sigma_{4}}(\bm{q}),$ (12)
+$\chi^{RPA}_{+-}(\bm{q})=\sum_{l_{1},l_{2}}\sum_{\{\sigma\}}\sigma^{+}_{\sigma_{1}\sigma_{2}}\sigma^{-}_{\sigma_{4}\sigma_{3}}\chi^{l_{1}\sigma_{1}l_{1}\sigma_{2}}_{l_{2}\sigma_{3}l_{2}\sigma_{4}}(\bm{q}).$
+The RPA spin susceptibilities corresponding to the Fermi surfaces shown in Fig.
+2 (a), (f), (k) and (p) can be calculated using Eqs.
+4, 6 and 12.
+The difference between longitudinal and transverse susceptibility originates from the breaking of spin-rotation invariance induced by spin-splitting effects such as RSOC here.
+Specifically, in the single-orbital limit, the longitudinal susceptibility, as illustrated in Fig.
+2(l), can be simplified as
+$\chi^{RPA}_{zz}(\bm{q})=\sum_{\mu\nu,k\sigma,l_{1}l_{2}}C_{l_{1},l_{2}}^{\mu\nu}[1-\cos(\theta_{k}-\theta_{k+q})]\times g[E],$ (13)
+where $e^{i\theta_{k}}=(\sin k_{y}+i\sin k_{x})/\sqrt{\sin^{2}k_{y}+\sin^{2}k_{x}}$, $g[E]=\frac{f(E_{\nu}(\bm{k}+\bm{q}))-f(E_{\mu}(\bm{k}))}{E_{\nu}(\bm{k}+\bm{q})-E_{\mu}(\bm{k})+i0^{+}}$ and the coefficient $C_{l_{1},l_{2}}^{\mu\nu}$ depend on $l_{1},l_{2},\mu,\nu$ and $k$ with $\mu$, $\nu$ denoting the band indices.
+For intra-pocket crossings, where k=(k+q)k=-(k+q), one has $\theta_{k}=\theta_{k+q}+\pi$, and therefore $1-\cos(\theta_{k}-\theta_{k+q})=2$, so that the susceptibility reaches its maximum.
+In contrast, for inter-pocket crossings this condition is not satisfied, and the susceptibility does not reach its maximum value.
+This explains why two distinct types of peaks are observed in Fig.
+2(l).
+Meanwhile, the transverse susceptibility, shown in Fig.
+2(m), can be written as
+$\chi^{RPA}_{+-}(\bm{q})=\sum_{\mu\nu,k}|a^{\uparrow}_{\mu}(k)|^{2}|a^{\downarrow}_{\nu}(k+q)|^{2}\times g[E].$ (14)
+The wave-function factor $|a^{\uparrow}_{\mu}(k)|^{2}|a^{\downarrow}_{\nu}(k+q)|^{2}$ is nearly constant (approximately 0.25) for both intra-pocket and inter-pocket crossings.
+As a result, three types of peaks appear, corresponding to intra-pocket crossing of the outer Fermi surface, inter-pocket crossing between different Fermi surfaces, and intra-pocket crossing of the inner Fermi surface.
+[…]
+We then examine the impact of transverse $\chi^{RPA}_{+-}(\bm{q})$.
+The numerical results indicate that for $\lambda_{so}=3$ (two-orbital limit) as shown in Fig.
+2 (h) and (r), $\chi^{RPA}_{+-}(\bm{q})$ is approximately zero, leading to a comparable strength of effective interactions between intra-pocket and inter-pocket interactions, as shown in Fig.
+2 (i) and (j) or (s) and (t).
+A stronger inter-pocket interaction favors a inter-pocket sign-change gap function making $\tau^{0}\sigma^{0}$ channel more likely to contribute to this sign change.
+Conversely, for $\lambda_{so}=0$ (one-orbital limit), the relatively strong $\chi^{RPA}_{+-}(\bm{q})$ depicted in Fig.
+2 (c) and (m) renders the effective inter-pocket interaction being significantly weaker than the intra-pocket interaction as shown in Fig.
+2 (d) and (e) or (n) and (o), thereby reducing the contribution of the $\tau^{0}\sigma^{0}$ channel to the inter-pocket sign-change behavior in the gap function.
+This indicates that the $\chi^{RPA}_{+-}(\bm{q})$ in Fig.
+2 (c), (h), (m) and (r) plays a important role in influencing the contributions from different interaction channels, resulting in the $\tau^{0}\sigma^{0}$ channel of the two-orbital model being more significant and stable than that of the model in single-orbital limit, thereby distinguishing it from the latter.
+IV Parity-mixed state and $S_{\pm}$-wave state
+Once superconductivity is induced by antiferromagnetic fluctuations, the pairing symmetry of the superconducting state can be classified according to the IRs of the $C_{4v}$ group.
 Table 1 lists the possible pairings with their form factors up to the next-nearest neighbor and their corresponding IRs.
-By solving the linear gap equation in Eq.
-(10), the leading superconducting pairing can be determined.
 Table 1: Possible IRs of superconducting pairings under the constraints of group $C_{4v}$. Here, $g_{k}=\sin k_{x}\sin k_{y}$.
@@
 $(\sin k_{x},-\sin k_{y},0)$ $B_{2}(B_{1})$
-$[(0,0,\sin k_{x}),(0,0,\sin k_{y})]$ $E$
+$(0,0,\sin k_{x}),(0,0,\sin k_{y})]$ $E$
 $i(\bm{d_{g}}\cdot\bm{\sigma})\sigma^{2}\tau^{2}$ $(0,0,\cos k_{x}+\cos k_{y})$ $A_{1}$
@@
 $\{(g_{k},0,0),(0,g_{k},0)\}$
-Figure 2: The phase diagrams of superconducting states as functions of some parameters.
-(a)-(b) $\lambda_{R}$-$\mu$ phase diagram.
-$U=4.5$ and $J=0.03U$ in (a).
-$U=5.4$ and $J=0.21U$ in (b).
-(c)-(d) $U$-$J$ phase diagram.
-$\lambda_{R}=0.9$ and $\mu=-5.3$ in (c).
-$\lambda_{R}=0.8$ and $\mu=-4.5$ in (d).
-In (a)-(d), all other parameters are the same as those in Fig.
-1 (b).
-In panel (a), the positions marked by numbers 1 and 2 correspond to the star-marked positions labeled 1 and 2 in (c) and (d).
-The notation $S_{\pm}$ and $d_{\pm}$ in $A_{1}$ and $B_{2}$ indicates that the superconducting gap functions on the two spin-split Fermi surfaces have opposite signs along the radial direction.
-The phase diagrams of the superconducting state with respect to $\lambda_{R}$-$\mu$ and $U$-$J$ are shown in Figs.
-2 and Figs.
-S4 and S6[43].
-Over a wide-range parameter space, we identify three possible leading pairing channels: $A_{1}$, $B_{1}$, and $B_{2}$.
-Notably, the superconducting states in the $A_{1}$ and $B_{2}$ channels exhibit highly unconventional sign-changing gap structures $A_{1}(S_{\pm})$ and $B_{2}(d_{\pm})$, which indicate that the superconducting gap functions on the two spin-split Fermi surfaces have opposite signs along the radial direction as shown in Fig.
-3 (a) and Fig.
-S9 (a)[43].
-Figure 3: (a)-(d) The superconducting gap functions projected onto the corresponding Fermi surfaces.
-The red and blue colors indicate the opposite signs of superconducting gap functions.
-(a) and (b) correspond to full and $A_{1,1}=i\sigma^{2}\tau^{0}\sigma^{0}\phi_{1k}$ channels in $A_{1}(S_{\pm})$ IR, respectively.
-(c) The renormalized gap function $\Delta_{A_{1}}(k_{F},\theta)/max(\Delta_{A_{1}}(k_{F},\theta))$ and $\Delta_{A_{1,1}}(k_{F},\theta)/max(\Delta_{A_{1,1}}(k_{F},\theta))$.
-(d) The weights of six sub-channels in $A_{1}(S_{\pm})$ IR as function as $t_{2}$.
-(e)-(f) The fitting gap function $\Delta_{1/2}(k)$ in Eq.
-13 along different high-symmetry lines $\Gamma-M$ and $\Gamma-X$.
-Other parameters in (a)-(d) are the same as those in Fig.
-1 (c) and of star labeled 2 in Fig.
-2 (a) and (d).
-To elucidate the origin of this unconventional sign-change superconducting state, we take the $A_{1}(S_{\pm})$ channel as an example.
-The gap function of noncentrosymmetric systems can be parameterized as follows[44],
-$\hat{\Delta}_{k}=[\hat{\phi}_{k}+\bm{\hat{d}_{k}}\cdot\bm{\sigma}]i\sigma^{2}.$ (12)
+By solving Eq.
+(10), the leading pairing can be determined.
+The phase diagrams concerning $\lambda_{R}-n$ with $n$ being the occupied particle number per site under various parameters are presented in Fig.
+3.
+Across a wide-range parameter space, we identify two possible leading pairing channels: $A_{1}(S_{\pm})$, and $B_{2}$ as shown in Figs.
+3 (a) and (b).
+Notably, the $A_{1}(S_{\pm})$ state exhibits a highly unconventional sign-change gap function, i.e., $S_{\pm}$-wave pairing, as shown in Figs.
+2 (a), (f), (k), (p).
+The phase diagrams in Fig.
+3 indicate that when the strength of RSOC is weak, the superconducting gap maintains a conventional $d$-wave gap, consistent with numerical results from the single-orbital model of cuprates.
+When the Hubbard $U$ is relatively small, a strong RSOC, leads to a inter-pocket sign-change superconducting gap as shown in Figs.
+3 (a) and (b).
+When $n$ is close to 0.5, the $A_{1}$ IR with an $s$-wave inter-pocket sign-change superconducting gap is dominant, while for a little smaller $n$, the $d$-wave inter-pocket sign-change gap labeled by $d_{\pm}$ become prominent.
+Additionally, when $n$ is around 0.2, there exists a region of $A_{1}$ IR with spin-triplet dominant.
+In the case where Hubbard $U$ is large shown in Fig.
+3 (c) and (d), spin density wave (SDW) states emerge in the large $n$ region, overshadowing the $A_{1}$ IR in large $n$ regime, resulting in a transition from spin-triplet-dominant to spin-singlet-dominant types in the $A_{1}$ gap function, as shown in Figs.
+3 (c) and (d).
+[…]
+Figure 4: (a)-(f) The plottings illustrate the contribution weights from six channels to the $A_{1}(S_{\pm})$-wave as functions as some specific parameters.
+The other fixed parameters are set as flollows: (a) $\lambda_{so}=3$, $n=0.26$ and $\lambda_{R}=0.35$, corresponding to Fig.
+2 (p), (b) $\lambda_{so}=3$, $n=0.26$ and $U=7.5$, (c) $\lambda_{R}=0.35$, $n=0.26$ and $U=7.5$, (d) $\lambda_{so}=3$, $n=0.5$ and $\lambda_{R}=0.9$, (e) $\lambda_{so}=3$, $n=0.5$ and $U=3.5$, (f) $\lambda_{R}=0.725$, $n=0.5$ and $U=3.5$.
+Note that in (a), (b), (d) and (e), when the condition requiring the absolute value of the weight is not imposed, the pairing in $\tau^{3}\sigma^{0}$ channel can acquire a minus sign under a direct projection.
+In (f), the right part of the green-line minimum is also taken in absolute value for the same reason.
+To investigate the parity-mixed states and elucidate the origin of this $S_{\pm}$-wave pairing, we express the gap function in a parameterized form as follows[46],
+$\hat{\Delta}_{k}=i\sigma^{2}[\hat{\phi}_{k}+\bm{\hat{d}_{k}}\cdot\bm{\sigma}].$ (15)
 In the one-orbital Rashba model, the projection of the gap function onto the two spin-split Fermi surfaces can be expressed in a simplified form, $\Delta_{1/2}(k)=\phi_{k}\pm|\hat{d}_{k}|$ with $\phi_{k}$ as a scalar.
 Thus, the aforementioned sign-change gap structure only occur when the triplet component is relatively strong and has helical $p$- or $f$-wave spin-triplet pairing, such as when the filling is near the van Hove singularity[32, 34, 35].
-Any doping away from the van Hove singularity consistently results in predominant d-wave spin-singlet pairing[33].
+Any doping away from the van Hove singularity consistently results in predominant nodal d-wave spin-singlet pairing[33].
 However, for the two-orbital model considered here, additional possibilities arise.
-We project the six sub-channels $i\sigma^{2}$($\tau^{0}\sigma^{0}\phi_{1k}$, $\tau^{3}\sigma^{0}\phi_{2k}$, $\tau^{0}\bm{\sigma}\cdot\bm{d_{3k}}$, $\tau^{3}\bm{\sigma}\cdot\bm{d_{4k}}$, $\tau^{1}\bm{\sigma}\cdot\bm{d_{5k}}$, $\tau^{2}\sigma^{3}d_{6k}$) in the spin-orbital representation, listed in Table 1 onto the two Fermi surfaces.
+We project the six sub-channels $i\sigma^{2}$$\Delta_{0}$($\tau^{0}\sigma^{0}\phi_{1k}$, $\tau^{3}\sigma^{0}\phi_{2k}$, $\tau^{0}\bm{\sigma}\cdot\bm{d_{3k}}$, $\tau^{3}\bm{\sigma}\cdot\bm{d_{4k}}$, $\tau^{1}\bm{\sigma}\cdot\bm{d_{5k}}$, $\tau^{2}\sigma^{3}d_{6k}$) in the spin-orbital representation, listed in Table 1 onto the two Fermi surfaces.
+The details can be found in Appendix B.
 Note that the first two sub-channels belong to spin-singlet while the latter four correspond to spin-triplet.
-The projection of the first sub-channel $i\sigma^{2}\tau^{0}\sigma^{0}\phi_{1k}$ is shown in Fig.
-3 (b), while the others are presented in Fig.
-S7 [43].
-We find that their superconducting gap functions exhibit the same projected distribution on the Fermi surfaces, ensuring consistency with the overall superconducting gap projection, as illustrated in Figs.
-3 (a), (b) and (c).
-[…]
-$\Delta_{1/2}(k)=\sqrt{2(1+\lvert B_{\bm{k}}^{\pm}\lvert^{2})}\phi_{1k},$ (13)
-with $B^{\pm}_{\bm{k}}=i(\sqrt{\lambda^{2}+h_{\bm{k}}^{2}\lambda_{R}^{2}}\pm\lambda_{R}h_{\bm{k}})/\lambda$ and $h_{\bm{k}}=\sqrt{\sin^{2}k_{x}+\sin^{2}k_{y}}$.
-In Fig.
-3 (e) and (f), we plot $\Delta_{1/2}(k)$ along the $\Gamma-X$ and $\Gamma-M$ directions, respectively.
-The sign-change feature of the gap function at the two Fermi momenta is clearly visible.
-Such sign-change gap structure is consistent with the Fermi surface nesting and $\chi^{zz}(\bm{q})$ in Figs.
-1 (c)-(e).
-Similar analysis can be performed for channel $B_{2}(d_{\pm})$, as shown in Figs.
-S8-S10[43].
-The results indicate that its sign-change feature cannot come from the pure spin-singlet channel and depends on the mixing of spin-singlet and spin-triplet components[43], in contrast to the $A_{1}(S_{\pm})$ pairing.
-Topological superconductivity— For noncentrosymmetric systems with time-reversal symmetry, the criteria for identifying TSC require that a pair of spin-split Fermi surfaces enclose an odd number of time-reversal invariant momentum points, and the order parameter must have opposite signs on the two Fermi surfaces[45].
+We further distinguish $A_{1}(S_{\pm})$ states according to the weights of each channel with varying of some parameters, as shown in Fig.
+4.
+This demonstrates that the superconducting state exhibits rich parity-mixed characteristics.
+Note that the emergence of $B_{2}(d_{\pm})$ state requires strong RSOC; an example with $\lambda_{R}=0.95$ is provided in Appendix C.
+Figures 4 clearly demonstrate that the $S_{\pm}$-wave state exhibits parity-mixed, with the relative contributions of the each channel depending on specific model parameters.
+This tunability makes our model a promising platform for exploring parity-mixed superconductivity, should it be realized in future material systems.
+As shown in Figs.
+4 (a) and (d), the ranking of the contributions from each channel is almost unaffected by Hubbard $U$; however, in certain cases where channel contributions are similar, Hubbard $U$ can alter the leading channel as shown in Fig.
+4 (a).
+For nn=0.26, fig.
+4 (b) illustrates that when $\lambda_{R}$ is small, the $S_{\pm}$-wave state is predominantly even-parity channels ($\tau^{0}\sigma^{0}$ and $\tau^{2}\sigma^{3}$).
+As $\lambda_{R}$ increases, their contribution diminishes, while the odd-parity components increase.
+This suggests that the spin-triplet pairing in this system is primarily induced by $\lambda_{R}$, fundamentally distinct from the intrinsic $p$-wave or $f$-wave triplet pairing mediated by ferromagnetic fluctuations near a van Hove singularity via doping.
+In contrast, our $S_{\pm}$-wave state is mediated by antiferromagnetic fluctuations, consistent with this mechanism.
+For nn=0.5 as shown in fig.
+4 (e), the dominant behavior of even-parity channels becomes even more pronounced, and over a wide range of the parameter $\lambda_{R}$, the spin-singlet channel remains the leading.
+This further indicates that the origin of the $S_{\pm}$ state is unrelated to $p$-wave or $f$-wave pairing.
+Therefore, in both weak and strong RSOC regime, the gap function of the $S_{\pm}$-wave state can be mainly governed by the singlet component, regardless of the presence of triplet contributions—even though the triplet channels can produce the same gap sign distribution, as shown in Fig.
+A3 in Appendix.
+Figure 4 (c) and (f) indicates that $\tau^{0}\sigma^{0}$ can be enhanced by $\lambda_{so}$.
+As discussed in relation to Fig.
+2, the introduction of $\lambda_{so}$ enhances the contribution of $\tau^{0}\sigma^{0}$ by canceling transverse susceptibility, ultimately strengthening inter-pocket interactions, which also enhances another even-parity channel, $\tau^{2}\sigma^{3}$.
+V Topological superconductivity
+For noncentrosymmetric systems with time-reversal symmetry, the criteria for identifying TSC require that a pair of spin-split Fermi surfaces enclose an odd number of time-reversal invariant momentum points such as $\Gamma$, $X$ and $M$ points, as shown in Fig.
+1 (b), where the Hamiltonian remains invariant under the time-reversal operation, and the order parameter must have opposite signs on the two Fermi surfaces[47].
 Additionally, the fully gapped superconducting state is preferable.
 According to this criterion, the Fermi surface in Fig.
-1 (c) and the corresponding superconducting gap in Fig.
-3 (a) or (b) support TSC.
-Figure 4: (a) The Wilson loop calculation for the $A_{1}(S_{\pm})$ superconducting state.
-(b) The quasi-particle spectrum of $A_{1}(S_{\pm})$ superconducting state.
+2 (p), with the $S_{\pm}$-wave pairing we focus on, support TSC.
+Figure 5: (a) The Wilson loop[48] calculation for the $S_{\pm}$ superconducting state.
+(b) The quasi-particle spectrum of $S_{\pm}$ superconducting state.
 Note that the in-gap edge states are double-degeneracy with opposite chriality.
@@
 In both (a) and (b), the parameters are the same as those in Fig.
-3 (a).
+2 (p).
+The parameter $\Delta_{0}$ represents the coefficient of the form factors $\phi_{i,k}$ and $\bm{d_{j,k}}$ in the superconducting gap function and is set to 0.1.
 For the two-dimensional system with time-reversal symmetry considered here, the Chern number is always zero, and the topological nature of the superconducting state can be characterized by a topological invariant $Z_{2}$.
-Due to the lack of spatial inversion symmetry, the $Z_{2}$ invariant can be determined by calculating the Wannier centers of the superconducting state’s wave functions[43, 46] and verified through the computation of the edge spectrum[47, 48].
+Due to the lack of spatial inversion symmetry, the $Z_{2}$ invariant can be determined by calculating the Wannier centers of the superconducting state’s wave functions[48] and verified through the computation of the edge spectrum[49, 50].
+The details can be found in Appendix B.
 Figs.
-4 (a) and (b) show the the Wannier center and edge spectrum, respectively, for the superconducting state with the Fermi surface from Fig.
-1 (c) and the gap structure from Fig.
-3 (b).
+5 (a) and (b) show the Wannier center and edge spectrum, respectively, for the superconducting state with the Fermi surface from Fig.
+2 (p) and the gap structure from Fig.
+A2.
 It is straightforward to conclude that this superconducting state has a topological invariant $Z_{2}=1$.
-The superconducting state in $B_{2}(d_{\pm})$ channel has nodes, however, it is also topologically nontrivial[43, 49], and exhibites Majorana flat bands and dispersive Majorana states on edges.
-Outlook and conclusion— The TSC state driven by predominant $s$-wave spin-singlet pairing in our study has significant implications.
+The $B_{2}(d_{\pm})$ state has nodes, however, it is also topologically nontrivial[51], and exhibites Majorana flat bands and dispersive Majorana states on edges (See Appendix C).
+VI Outlook and conclusion.
+The TSC state driven by predominant spin-singlet pairing in our study has significant implications.
 First, TSC states realized through $p$- or $f$-wave spin-triplet pairing are extremely rare in real material systems.
-In contrast, almost all reported superconductors exhibit $s$-wave spin-singlet pairing, including cuprate, iron-based, and nickelate high-temperature superconductors[50, 51, 52].
+In contrast, almost all reported superconductors exhibit spin-singlet pairing, including cuprate, iron-based, and nickelate high-temperature superconductors[52, 53, 54].
 Second, the splitting of Fermi surfaces accompanied by spin-texture characteristics due to broken inversion symmetry has been experimentally reported in systems such as cuprate and iron-based superconductors[30, 31].
@@
 In summary, we investigated superconductivity in a new class of correlated non-centrosymmetric systems.
-Using the theory of antiferromagnetic-fluctuation-mediated superconductivity, we identified four possible superconducting states $A_{1}(S_{\pm})$, $B_{1}$, $B_{2}$ and $B_{2}(d_{\pm})$ under different model parameters.
-all of which are predominantly driven by spin-singlet pairing with $A_{1}$, $B_{1}$ and $B_{2}$ pairing symmetries.
-Notably, the superconducting state with $A_{1}(S_{\pm})$ IR is predominantly driven by spin-singlet pairing and belongs to a bulk topological superconducting state characterized by a $Z_{2}=1$ topological invariant.
+Using the theory of antiferromagnetic-fluctuation-mediated superconductivity, we identified three possible superconducting states $A_{1}(S_{\pm})$, $B_{2}$ and $B_{2}(d_{\pm})$ under different model parameters.
+Notably, all states are parity-mixed, and the ratio between even-parity and odd-parity components can be tuned by the model parameters.
+More importantly, the $A_{1}(S_{\pm})$ state belongs to a bulk topological superconducting state characterized by a $Z_{2}=1$ topological invariant.
+In both weak and strong RSOC regime, the $A_{1}(S_{\pm})$ state can be predominantly driven by spin-singlet components, while $B_{2}(d_{\pm})$ is a nodal TSC and only emerges in strong RSOC regime.
 Our proposal is different from all present scenarios based on proximity effects and strong $p$- or $f$-wave spin-triplet pairing.
-We further highlighted the potential realization of such topological superconducting states in certain high-temperature superconductors, providing a new pathway for exploring next-generation high-temperature topological superconductors.
 Acknowledgements.
 This work was financially supported by the National Key R&D Program of China (Grants No.
-2022YFA1403200, No.
-2024YFA1613200 and Grant No.
+2024YFA1613200, No.
+2022YFA1403200 and Grant No.
 2023YFA1407300), National Natural Science Foundation of China (Grants No.
@@
 AHHM-FX-2020-02.
+Appendix A Formalism of the multi-orbital RPA approach
+We calculate the susceptibility using the expression
+$\chi^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{q},i\nu_{n})=\sum_{\bm{k},\bm{k}^{\prime}}\int_{0}^{\beta}d\tau\braket{T_{\tau}c^{\dagger}_{l_{1}\sigma_{1}}(\bm{k}+\bm{q},\tau)c_{l_{2}\sigma_{2}}(\bm{k},)c^{\dagger}_{l_{4}\sigma_{4}}(\bm{k}^{\prime}-\bm{q},0)c_{l_{3}\sigma_{3}}(\bm{k}^{\prime},0)}$ (16)
+The Green’s function expression of the lowest order is given by
+$\chi^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{q},i\nu_{n})=-T\sum_{\bm{k},i\omega_{n}}G_{l_{3}\sigma_{3}l_{1}\sigma_{1}}(\bm{k},i\omega_{n})G_{l_{2}\sigma_{2}l_{4}\sigma_{4}}(\bm{k}+\bm{q},i\omega_{n}+i\nu_{n}),$ (17)
+$\chi^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{q},i\nu_{n})=-\frac{1}{N}\sum_{\bm{k},\mu\nu}a^{l_{3}\sigma_{3}}_{\mu}(\bm{k})a^{l_{1}\sigma_{1}*}_{\mu}(\bm{k})a^{l_{2}\sigma_{2}}_{\nu}(\bm{k}+\bm{q})a^{l_{4}\sigma_{4}*}_{\nu}(\bm{k}+\bm{q})\frac{f(E_{\nu}(\bm{k}+\bm{q}))-f(E_{\mu}(\bm{k}))}{E_{\nu}(\bm{k}+\bm{q})-E_{\mu}(\bm{k})+i\nu_{n}}$ (18)
+where the Matsubara Green’s function is
+$\hat{G}(\bm{k},i\omega_{n})=[i\omega_{n}-H_{0}(\bm{k})]^{-1}.$ (19)
+Here, the indices $l_{1-4}$ refer to orbital indices, while the indices $\sigma_{1-4}$ refer to spin indices.
+$i\omega_{n}$ and $i\nu_{n}$ denote Fermionic and Bosonic Matsubara frequencies, respectively.
+In practical calculations, we can express the susceptibility as Eq.
+(18) where $N$ is the number of $\bm{k}$-points in the Brillouin Zone, $a^{l\sigma}_{\mu}(\bm{k})=\braket{l\sigma,\bm{k}|\mu,\bm{k}}$ is the wave function in energy space, and $f(E)$ is the Fermi-Dirac distribution.
+For the calculation of static susceptibility, we set $i\nu_{n}=i0^{+}$ on the right side of equation (18), while on the left side, we have $\chi^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{q},i\nu_{n}=0)$ to smooth results of the calculation.
+The Hubbard interaction matrix is given by
+$W^{l_{1}\sigma l_{1}\bar{\sigma}}_{l_{1}\sigma l_{1}\bar{\sigma}}=U$ $W^{l_{1}\sigma l_{1}\sigma}_{l_{1}\bar{\sigma}l_{1}\bar{\sigma}}=-U$ (20)
+$W^{l_{2}\sigma l_{1}\bar{\sigma}}_{l_{2}\sigma l_{1}\bar{\sigma}}=V$ $W^{l_{1}\sigma l_{1}\sigma}_{l_{2}\bar{\sigma}l_{2}\bar{\sigma}}=-V$
+$W^{l_{1}\sigma l_{1}\bar{\sigma}}_{l_{2}\sigma l_{2}\bar{\sigma}}=J$ $W^{l_{1}\sigma l_{2}\sigma}_{l_{1}\bar{\sigma}l_{2}\bar{\sigma}}=-J$
+$W^{l_{1}\sigma l_{2}\bar{\sigma}}_{l_{1}\sigma l_{2}\bar{\sigma}}=J^{\prime}$ $W^{l_{1}\sigma l_{2}\sigma}_{l_{2}\bar{\sigma}l_{1}\bar{\sigma}}=-J^{\prime}$
+with all other matrix elements equal to zero. The RPA susceptibility is given by
+$\chi^{RPA}(\bm{q})=[1-\chi(\bm{q})W]^{-1}\chi(\bm{q}),$
+where the charge susceptibility is
+$\chi^{RPA}_{c}(\bm{q})=\sum_{l_{1},l_{2}}\sum_{\{\sigma\}}\sigma^{0}_{\sigma_{1}\sigma_{2}}\sigma^{0}_{\sigma_{4}\sigma_{3}}\chi^{l_{1}\sigma_{1}l_{1}\sigma_{2}}_{l_{2}\sigma_{3}l_{2}\sigma_{4}}(\bm{q}),$
+the longitudinal susceptibility is
+$\chi^{RPA}_{zz}(\bm{q})=\sum_{l_{1},l_{2}}\sum_{\{\sigma\}}\sigma^{z}_{\sigma_{1}\sigma_{2}}\sigma^{z}_{\sigma_{4}\sigma_{3}}\chi^{l_{1}\sigma_{1}l_{1}\sigma_{2}}_{l_{2}\sigma_{3}l_{2}\sigma_{4}}(\bm{q}),$
+and the transverse susceptibility is
+$\chi^{RPA}_{+-}(\bm{q})=\sum_{l_{1},l_{2}}\sum_{\{\sigma\}}\sigma^{+}_{\sigma_{1}\sigma_{2}}\sigma^{-}_{\sigma_{4}\sigma_{3}}\chi^{l_{1}\sigma_{1}l_{1}\sigma_{2}}_{l_{2}\sigma_{3}l_{2}\sigma_{4}}(\bm{q}),$
+After applying the RPA, we are able to calculate the effective interaction of Cooper pairs in orbital-spin space.
+The relevant Feynman diagrams include ladder and bubble diagrams, which are illustrated in FIG.
+A1.
+The total effective pairing interaction can be expressed as
+Figure A1: The second-order bubble and ladder diagram
+$V^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{k},\bm{k^{\prime}})=W^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}+V_{l}(\bm{k},\bm{k^{\prime}})-V_{b}(\bm{k},\bm{k^{\prime}}),$ (21)
+with the ladder term in FIG. A1 (b)
+$V_{l}(\bm{k},\bm{k^{\prime}})=[\hat{W}\hat{\chi}^{RPA}\hat{W}]^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{k}+\bm{k^{\prime}}),$ (22)
+and the bubble term in FIG. A1 (a)
+$V_{b}(\bm{k},\bm{k^{\prime}})=[\hat{W}\hat{\chi}^{RPA}\hat{W}]^{l_{1}\sigma_{1}l_{4}\sigma_{4}}_{l_{3}\sigma_{3}l_{2}\sigma_{2}}(\bm{k}-\bm{k^{\prime}}).$ (23)
+The potential superconducting pairing can be evaluated by solving the following linearized Eliashberg equation,
+$\lambda\Delta_{l_{1}\sigma_{1}l_{4}\sigma_{4}}(\bm{k})=T\sum_{\bm{k^{\prime}}i\omega_{n}}\sum_{l_{2}l_{3}\sigma_{2}\sigma_{3}}V^{l_{1}\sigma_{1}l_{2}\sigma_{2}}_{l_{3}\sigma_{3}l_{4}\sigma_{4}}(\bm{k},\bm{k^{\prime}})F_{l_{3}\sigma_{3}l_{2}\sigma_{2}}(\bm{k^{\prime}},i\omega_{n})$ (24)
+with
+$F_{l_{3}\sigma_{3}l_{2}\sigma_{2}}(\bm{k^{\prime}},i\omega_{n})=G_{l_{3}\sigma_{3}l\sigma}(\bm{k^{\prime}},i\omega_{n})\Delta_{l\sigma l^{\prime}\sigma^{\prime}}(\bm{k^{\prime}})G_{l_{2}\sigma_{2}l^{\prime}\sigma^{\prime}}(-\bm{k^{\prime}},-i\omega_{n}).$ (25)
+Here, $\Delta_{l_{1}\sigma_{1}l_{4}\sigma_{4}}(\bm{k})$ is the superconducting pairing function, and the leading superconducting instability can be identified by the largest eigenvalue $\lambda_{\alpha}$.
+Appendix B BdG Hamiltonian projection between spin-orbital space and band space, and calculation of $Z_{2}$
+We now focus on the $S_{\pm}$-wave states in orbital-spin space. The BdG Hamiltonian is given by
+$H_{BdG}=\left[\begin{aligned} H_{0}(k)&&\Delta(k)\\ \Delta^{*}(k)&&-H^{*}_{0}(-k)\end{aligned}\right].$ (26)
+Here, the normal part $H_{0}(k)$ can be can be diagonalized using $\hat{V}^{-1}(k)H_{0}(k)\hat{V}(k)=\hat{E}(k)$, with $\hat{E}_{\mu\nu}=E_{\mu}(k)\delta_{\mu\nu}$.
+$\mu$ and $\nu$ are the band indices.
+$\hat{V}(k)$ can be parameterized as follows:
+$\hat{V}(k)=\left[\begin{aligned} C_{1}e^{i\theta}&&-C_{3}e^{i\theta}&&C_{4}e^{i\theta}&&C_{2}e^{i\theta}\\ -iC_{1}&&-iC_{3}&&iC_{4}&&iC_{2}\\ -iC_{2}e^{i\theta}&&iC_{4}e^{i\theta}&&iC_{3}e^{i\theta}&&-iC_{1}e^{i\theta}\\ C_{2}&&C_{4}&&C_{3}&&C_{1}\end{aligned}\right],$ (27)
+where $e^{i\theta}=(\sin k_{y}+i\sin k_{x})/h_{\bm{k}}$, $2(C_{1}^{2}+C_{2}^{2})=2(C_{3}^{2}+C_{4}^{2})=1$ and $h_{\bm{k}}=\sqrt{\sin^{2}k_{x}+\sin^{2}k_{y}}$.
+For the pairing function belonging to the $A_{1}$ irreducible representation, it can be rewritten as
+$\Delta(k)/(i\sigma^{2}\Delta_{0})$ $=\tau^{0}\sigma^{0}\phi_{1k}+\tau^{0}\bm{\sigma}\cdot\bm{d_{3k}}+\tau^{2}\sigma^{3}d_{6k}+\tau^{1}\bm{\sigma}\cdot\bm{d_{5k}}+\tau^{3}\sigma^{0}\phi_{2k}+\tau^{3}\bm{\sigma}\cdot\bm{d_{4k}}$ (28)
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Here, $\phi_{1k}$, $\phi_{2k}$ and $d_{6k}$ has general form of extended s-wave gap function, $\bm{d_{3k}}=\bm{d_{4k}}\propto(-sink_{y},sink_{x},0)$ and $\bm{d_{5k}}\propto(sink_{x},sink_{y},0)$.
+By solving the Eliashberg function in Eq.
+(24), the numerical results in orbital-spin space can be obtained and are presented in Fig.
+A2.
+The fitting results as listed as Eq.
+(29)
+Figure A2: Gap function in six possible channel of $A_{1}$ IR under the condition of Fig.
+2 (p) over the angle $\theta$ on the two Fermi surfaces.
+Here, the defination of polar angle $\theta$ can be found in Fig.
+A3 (a).
+$\Delta_{0}$ is a constant determined by a cut-off energy.
+Figure A3: (a) The projected gap function of full $A_{1}$ IR in band space, normalized to its maximum.
+Here, the polar angle $\theta$ is defined in the polar coordinate system.
+(b)-(g) The projected gap function of six individual sub-channels of $A_{1}$ IR in band space, renormalized by the maximum of $i\sigma^{2}\tau^{0}\sigma^{0}\phi_{1k}$.
+Here U=7.5 and J=0.
+Other parameters are the same with Fig.
+2 (p) in main text.
+Do the projection, the effective gap function in band space of $H_{0}$ is given by
+$\Delta_{1/2}(k)=\Delta_{0}(\phi_{1k}+2(C_{1/3}^{2}-C_{2/4}^{2})\phi_{2k}-4C_{1/3}C_{2/4}d_{6k}\mp C^{t}_{1/2})$ (30)
+where
+$C^{t}_{1/2}=2(C_{1/3}^{2}-C_{2/4}^{2})|d_{3k}|-4C_{1/3}C_{2/4}|d_{5k}|+|d_{4k}|$ (31)
+in this case, $C_{2}/C_{1}<<1$, $C_{4}/C_{3}<<1$, then we have
+$\Delta_{1/2}(k)\approx\Delta_{0}(\phi_{1k}+\phi_{2k}-2C_{2/4}d_{6k}\mp C^{t}_{1/2})$ (32)
+where
+$C^{t}_{1/2}\approx|d_{3k}|-2C_{2/4}|d_{5k}|+|d_{4k}|$ (33)
+Note that the gap functions in Eq.
+28 and Eq.
+32 are expressed in spin-orbital and band basis, respectively.
+They exhibit a one-to-one correspondence.
+In Fig.
+A3 (a), we present the distribution of the total projected gap function on the Fermi surface in the band basis, showing a clear $S_{\pm}$-wave characteristic.
+In Figs.
+A3 b–g, we show the distributions of the projected gap function on the Fermi surface in the band basis when each subchannel in the spin-orbital basis is individually nonzero.
+It can be seen that the two spin-singlet subchannels carry the strongest weight, and all subchannels share the same $S_{\pm}$-wave characteristic.
+Note that for the remaining two bands, which lie far from the Fermi level, the gap function acquires an opposite sign; however, these contributions can be safely ignored since superconductivity is governed primarily by the low-energy states near the Fermi surface.
+Given that the system is time-reversal invariant, the Chern number is always zero.
+Therefore we have to use the $Z_{2}$ number to study its topological properties.
+We employ Wilson’s loop approach [55] where the non-Abelian Berry connection is defined as
+$F^{mn}_{i,i+1}=\braket{n,k_{x,i},k_{y}|m,k_{x,i+1},k_{y}}.$ (34)
+Multiplying $F^{mn}_{i,i+1}$ from A to Z including $F^{mn}_{N,1}$, a new matirx $D(k_{y})$ can be obtained as follow,
+$D(k_{y})=F_{1,2}F_{2,3}...F_{N,1}.$ (35)
+Here, $D(k_{y})$ is a $4\times 4$ matrix whose eigenvalue phases $\theta^{D}_{m}(k_{y})$ represent the evolution of the Wannier center pair in $k_{y}\in[-\pi,\pi]$.
+The integer winding number of the Wannier center pair can be viewed as the number of times it encloses the original point.
+Figure A4: The distribution of the gap functions for the six subchannels in the $d_{\pm}$-wave states as a function of the angle $\theta$ on the two Fermi surfaces is presented.
+The parameters used are $\lambda_{R}=0.95$, $\lambda_{so}=3$, $n=0.26$, $U=3.5$ and other parameters remain consistent with main text.
+Figure A5: (a) The projected gap function of full $B_{2}$ IR in band space, normalized to its maximum.
+(b)-(g) The projected gap function of six individual sub-channels of $B_{2}$ IR in band space, renormalized by the maximum of $i\sigma^{2}\tau^{0}\sigma^{0}\phi_{1k}$.
+Here parameters are the same with Fig.
+A4.
+Figure A6: (a) The quasi-particle spectrum of $d_{\pm}$ superconducting state in (100) direction. (b) The quasi-particle spectrum of $d_{\pm}$ superconducting state in (110) direction.
+Appendix C Properties of $d_{\pm}$ waves
+An addtional noteworthy class of superconducting gaps identified in our calculations is the $B_{2}(d_{\pm})$ wave.
+To obtain this result, we increased the strength of the RSOC to $\lambda_{R}=0.95$, surpassing the onsite SOC of $\lambda_{so}=3$.
+Figure (A4) clearly demonstrates that the wave state exhibits parity mixing, indicating that the sign changes across different Fermi surfaces are not solely attributed to spin-triplet subchannels; rather, spin-singlet subchannels inherently exhibit sign changes as well.
+In the spin-orbital representation, the superconducting gap function can be fitted up to 6th-nearest-neighbor terms.
+Specifically, $\phi_{1k}=0.138\sin k_{x}\sin k_{y}-0.005\sin 2k_{x}\sin 2k_{y}-0.05(\sin 2k_{x}\sin k_{y}+\sin k_{x}\sin 2k_{y})$.
+In Fig.
+A5 (a), we present the distribution of the total projected gap function on the Fermi surface in the band basis, showing a clear $d_{\pm}$-wave characteristic.
+In Figs.
+A5 b–g, we show the distributions of the projected gap function on the Fermi surface in the band basis.
+It can be seen that the two spin-singlet subchannels is naturely sign changed and carry strong weight, and all subchannels share the same $d_{\pm}$-wave characteristic.
+Regarding the spectrum for a boundary that breaks translational periodicity along the (100) direction, two types of edge states appear.
+The first is a conventional flat-band edge state, analogous to that found in standard $B_{2}$ pairing.
+The second is a pair of Dirac-type edge states, which arise from the sign change of the superconducting gap between different Fermi surfaces, as shown in Fig.
+A6(a).
+It is important to note that these Dirac-type edge states may not always be directly observable, as they can be obscured by bulk states.
+However, when the boundary is oriented along an alternative direction, such as (110), these edge states can always be detected, since there is no node at $k_{(110)}=0$ as demonstrated in Fig.
+A6(b).